# Masked Audio Generation using a Single Non-Autoregressive Transformer

### 09 Jan 2024 v1

@@ 4 Experimental setup
Evaluation metrics. We evaluate the proposed method using the same setup as proposed by Copet et al. (2023); Kreuk et al. (2022a), which consists of both objective and subjective metrics. For the objective metrics, we use: the Fréchet Audio Distance (FAD), the Kullback-Leiber Divergence (KL), and the CLAP score (CLAP). We report the FAD (Kilgour et al., 2018) using the official implementation in Tensorflow with the VGGish model 222github.com/google-research/google-research/tree/master/frechet_audio_distance. Following Kreuk et al. (2022a), we use a state-of-the-art audio classifier trained for classification on AudioSet (Koutini et al., 2021) to compute the KL-divergence over the probabilities of the labels between the original and the generated audio. We also report the CLAP score (Wu et al., 2023; Huang et al., 2023b) between the track description and the generated audio to quantify audio-text alignment, using the official CLAP model  [^3] https://github.com/LAION-AI/CLAP.
@@ 5.1 Text-to-music generation
Figure 2: Latency and throughput analysis: MAGNeT is particularly suited to small batch sizes (its latency is up to 10 times lower than MusicGen), while MusicGen benefits from a higher throughput for bigger batch sizes. MAGNeT offers flexibility regarding the latency/quality trade off by allowing a customizable decoding schedule or using an autoregressive-generated prompt of flexible duration in the Hybrid-MAGNeT variant.
@@ 5.2 Analysis
Compared with the baseline autoregressive model (red curve), the non-autoregressive model (dashed blue curve) especially excels on small batch sizes due to parallel decoding, with a latency as low as $600$ ms for a single generated sample (batch size of two in the Fig. 1(a)), more than $10$ times faster than the autoregressive baseline. This is especially interesting in interactive applications that require low-latency processing. The non-autoregressive model keeps being faster than the baseline until a batch size of $64$ is reached.
@@ 5.3 Ablation
Table 3: Rescorer ablation. We evaluate the effect of the rescorer on model performance. For human studies, we report mean and CI95.
@@ Appendix D Model inference
5%**** 08_appendix.tex Line 125 ****gen_seq = torch.full((B, T), mask_id, dtype=torch.long)

### 05 Mar 2024 v2

@@ 4 Experimental setup
Evaluation metrics. We evaluate the proposed method using the same setup as proposed by Copet et al. (2023); Kreuk et al. (2022a), which consists of both objective and subjective metrics. For the objective metrics, we use: the Fréchet Audio Distance (FAD), the Kullback-Leiber Divergence (KL), and the CLAP score (CLAP). We report the FAD (Kilgour et al., 2018) using the official implementation in Tensorflow with the VGGish model 222github.com/google-research/google-research/tree/master/frechet_audio_distance. Following Kreuk et al. (2022a), we use a state-of-the-art audio classifier (Koutini et al., 2021) to compute the KL-divergence over the probabilities of the labels between the original and the generated audio. We also report the CLAP score (Wu et al., 2023; Huang et al., 2023b) between the track description and the generated audio to quantify audio-text alignment, using the official CLAP model  [^3] https://github.com/LAION-AI/CLAP.
@@ 5.1 Text-to-music generation
Figure 2: Latency and throughput analysis: MAGNeT is particularly suited to small batch sizes (up to 10 times lower latency than MusicGen), while MusicGen benefits from a higher throughput for bigger batch sizes. MAGNeT offers flexibility regarding the latency/quality trade off by allowing a customizable decoding schedule or following the Hybrid-MAGNeT variant.
@@ 5.2 Analysis
Compared with the baseline autoregressive model (red curve), the non-autoregressive model (dashed blue curve) especially excels on small batch sizes due to parallel decoding, with a latency as low as $600$ ms for a single generated sample (batch size of two in the 1(a)), more than $10$ times faster than the autoregressive baseline. This is especially interesting in interactive applications that require low-latency. The non-autoregressive model is faster than the baseline up to a batch size of $64$.
@@ 5.3 Ablation
Table 3: We evaluate the effect of the rescorer on model performance. We report mean and CI95.
@@ 7 Discussion
Acknowledgements.
The authors would like to thank Or Tal, Michael Hassid and Nitay Arcusin for the useful theoretical discussions. The authors would additionally like to thank Kamila Benzina and Nisha Deo for supporting this project. This research work was supported in part by ISF grant 2049/22.
@@ Appendix D Model inference
5%**** 09_appendix.tex Line 125 ****gen_seq = torch.full((B, T), mask_id, dtype=torch.long)